1. Order: Amoxicillin 500 mg po q12hr .child weighs 40 lbs. safe range is 20-30 mg/kg/24hr

a. what is the safe 24 hours dose in mg?

b. is the ordered dose safe for a 24 hours period?

c. How many mg/ dose will the patient receive?

2 Order: phenytoin 60 mg po q 6 hr. Child weighs 57 lbs. The safe range is 8-10 mg/kg/24hr.

a. What is the safe 24 hr dose in mg

b Is the ordered dose safe for a 24 hr period

c If the medication is supplied 125 mg/5ml, how many ml will you administer per dose

3. Carafate 6 mg/kg/day q6hr is ordered for a child weighing 35 kg Carafate suspension is available in 50 mg/ml

a. How many mg per dose will the child receive?

b. How many mg per day will the child receive?

c. How many ml per dose will the child receive?
Answer #1
  1. Ordered Amoxicillin= 500 mg po q12hr

Child weight= 40 lbs= 18.18 kg

The safe range is =20-30 mg/kg/24hr

  1. The safe 24 hours dose in mg in this case will be=18.18X 20 to 18.18X30

                                                                              = 363.63mg - 545.4 mg

b. The ordered dose 500 mg po q12hr or 1000mg for 24 hrs is not safe as it is above the safe 24 hours dose rage i.e. 363.63mg - 545.4 mg

c. The patient will receive 500 mg per dose

2. Ordered phenytoin= 60 mg po q 6 hr.

Child weight= 57 lbs=25.90 kg

The safe range is 8-10 mg/kg/24hr.

a. The safe 24 hr dose in mg will be=25.90X8 mg to 25.90X 10mg= 207.27mg -259 mg/24 hors

b The ordered dose is =60 mg po q 6 hr. or 240 mg for 24 hours is safe as it is in the between the range of safe dose i.e. 207.27mg -259 mg/24

c The medication is supplied as 125 mg/5ml= 25 mg/ml, then the required ml per dose will be=?

60mg = 60/25 ml=2.4 ml

3.Ordered Carafate= 6 mg/kg/day q6hr

Child weighing 35 kg

Carafate suspension is available in 50 mg/ml

  1. Per dose the child will receive=35X6=210 mg/day q6hr

per dose =210/4=51.4 mg

b. The child will receive= 210 mg per day

c. How many ml per dose will the child receive= 51.5/50=1.08 ml per dose